# Learning Streaming Video Representation via Multitask Training

### v1

@@ Abstract
Understanding continuous video streams plays a fundamental role in real-time applications including embodied AI and autonomous driving. Unlike offline video understanding, streaming video understanding requires the ability to process video streams frame by frame, preserve historical information, and make low-latency decisions. To address these challenges, our main contributions are three-fold. (i) We develop a novel streaming video backbone, termed as StreamFormer, by incorporating causal temporal attention into a pre-trained vision transformer. This enables efficient streaming video processing while maintaining image representation capability. (ii) To train StreamFormer, we propose to unify diverse spatial-temporal video understanding tasks within a multitask visual-language alignment framework. Hence, StreamFormer learns global semantics, temporal dynamics, and fine-grained spatial relationships simultaneously. (iii) We conduct extensive experiments on online action detection, online video instance segmentation, and video question answering. StreamFormer achieves competitive results while maintaining efficiency, demonstrating its potential for real-time applications.
@@ 1 Introduction
Streaming video understanding poses several unique challenges compared to traditional offline approaches. First, efficient frame-by-frame processing is crucial, as incoming visual signals must be handled promptly and continuously. Unlike existing offline video backbones [33, 3, 98], which typically process clips as their basic unit, streaming models must operate seamlessly at a per-frame basis. Second, long-term context preservation is essential, requiring models to effectively capture temporal dependencies across frames under strict memory and computational constraints. Third, low-latency decision-making is indispensable, ensuring predictions or actions are promptly delivered for real-time scenarios. Addressing these challenges necessitates specialized backbone architectures specifically designed to balance efficiency, memory management, and accuracy.
@@
Recent works [119, 95] adopt pre-trained image encoders to extract frame features, which are subsequently processed auto-regressively by large language models (LLMs). Though this strategy aligns well with streaming video requirements, the reliance on aggregating coarse historical information within LLMs often leads to the loss of important fine-grained spatial and temporal details. To address these limitations, we propose training a streaming video backbone capable of supporting diverse spatiotemporal tasks, ranging from global scene understanding to fine-grained spatiotemporal event recognition. The proposed model, termed StreamFormer, adapts Vision Transformer architectures pre-trained on images (e.g. SigLIP [114]) by introducing causal attention along the temporal dimension at the spatial feature extraction stage. This approach allows the backbone to explicitly model temporal dependencies at early vision encoding stages, resulting in richer and more temporally coherent feature representations.
@@
To effectively train the streaming video backbone, we leverage the datasets that are manually annotated by the vision community [12, 36, 103, 11, 2, 35, 108], unifying various spatial-temporal video understanding tasks into a coherent multi-task learning framework. In particular, we combine three complementary learning objectives: (i) video-level classification for global semantics; (ii) per-frame supervision for temporal understanding, and (iii) pixel-wise segmentation for fine-grained spatial comprehension. By leveraging shared textual embeddings from SigLIP, our framework supports both closed-set and open-set tasks. This text-visual alignment mechanism facilitates knowledge transfer across tasks with varying granularities, enabling compatibility with diverse label spaces. In contrast to existing video backbone training paradigms [98, 122, 58] that predominantly rely on global contrastive learning using massive video-text datasets [99, 6, 67], our joint optimization of diverse objectives explicitly ensures capturing detailed temporal motion cues and spatial precision within streaming contexts.
@@
In summary, our contributions are threefold: (i) we propose StreamFormer, that effectively converts pre-trained image backbones into streaming video models via temporal causal attention and spatial low-rank adaptations. Our approach achieves efficient streaming video processing without sacrificing spatial representational quality; (ii) we introduce a hierarchical multitask learning paradigm, systematically unifying global-level, temporal-level, and spatial-level learning objectives. This paradigm enables our model to simultaneously comprehend global semantics, temporal dynamics, and fine-grained spatial relationships within streaming video; (iii) we empirically demonstrate the effectiveness and efficiency of our method on diverse downstream tasks, including online action detection, online video instance segmentation, and video question answering. Our approach achieves promising performance across these tasks, effectively balancing accuracy and computational efficiency in streaming scenarios.
@@ 2 Related Work
Video Backbone. Early video understanding methods employ 3D convolutional neural networks [12, 87, 32] or factorized convolutions [93, 88, 89] to learn spatial-temporal features. Recent works have adopted Vision Transformers [27] to capture long-range dependencies in videos by incorporating global spatiotemporal self-attention [3], local self-attention [65], local-global self-attention [51], divided space-time attention [7], and multi-scale attention mechanisms [31, 54]. Despite the progress in offline video tasks, the use of bi-directional attention in visual backbones makes them less suitable for real-time video tasks. In NLP, LLMs typically adopt a decoder-only architecture [85, 86, 59, 4], where causal attention enables efficient autoregressive generation during inference. Currently, autoregressive modeling in vision primarily focuses on training backbones for image or video generation [90, 72, 5, 81, 83]. For visual understanding tasks, iGPT [17] learns visual representations by predicting next pixels, while AIM [30] explores the scaling laws for autoregressive image pre-training. Toto [77] trains a causal video transformer to predict the next visual patch. Unlike these works, we target at designing StreamFormer with causal temporal attention to address various video understanding tasks.
@@
Video Language Pre-training. Building on the success of image-text pretraining [74, 114, 38], recent video-language foundation models [98, 100, 122, 106, 1] learn generalizable representations by training on massive video-text pairs [6, 99, 113]. UMT [50] and InternVideo2 [100] combine feature distilling, masked video modeling and video-text contrastive learning, to learn unimodal and multimodal representations. VideoPrism [122] integrates global-local distillation and token shuffling to enhance video representation learning. Despite strong performance, these methods typically require extensive training on million-scale video-text pairs. In contrast, we explore leveraging the rich, fine-grained human annotations available in existing video datasets for video language training by adopting a multitask learning paradigm, achieving higher data efficiency.
@@
Streaming Video Understanding. Streaming video understanding tasks necessitate the processing of video inputs on a frame-by-frame basis and demand timely reactions. Existing streaming video approaches mainly target at online action detection [104, 105, 15, 92], online video instance segmentation [101, 118, 111], dense video captioning [126]. Leveraging the autoregressive paradigm in LLMs, recent approaches extract frame-wise features using pre-trained image encoders and pass them to LLMs for online question answering [25, 116, 43]. However, aggregating coarse frame features often results in the loss of fine-grained spatial and temporal details. In contrast, StreamFormer addresses causal temporal interaction in the video backbone to preserve spatial-temporal details.
@@
Multitask Learning. Multitask models provide different outputs for a given input, simultaneously supporting various tasks [13, 79, 112, 29, 68, 46]. Motivated by the success of unifying diverse tasks as sequence-to-sequence modeling in LLMs [76, 9, 24], recent works either train a unified multimodal and multitask transformer model [115, 94, 66, 82, 127, 18], or leverage LLM as a bridge to combine task-specific heads [96, 102, 107]. Inspired by these works, we train our streaming video backbone in a unified multitask visual-language alignment framework.
@@
Figure 2: Overall framework of StreamFormer. Left: Our StreamFormer is trained under a unified visual-language alignment framework, enabling simultaneous understanding of global semantics, temporal dynamics, and fine-grained spatial relationships. Right: Each level utilizes features of different granularities: (i) last frame for the global level, (ii) per frame for the temporal level, (iii) and per frame per patch feature for the spatial level.
@@ 3.1 Problem Formulation
In this paper, we consider the task of streaming video understanding, that aims to continuously process video frames by extracting spatiotemporal features in real-time, while retaining historical context and supporting diverse video understanding tasks. To achieve this, we propose to train a unified streaming video backbone capable of capturing global semantics, temporal dependencies, and fine-grained spatial details, all within a single framework.
Formally, given a streaming video, represented as $\mathcal{V}^{T}=[I_{1},I_{2},\dots,I_{T}]$, where $T$ denotes the total number of frames in the video. At any time step $t\ (1\leq t\leq T)$, the visual features of different granularities can be computed as:
@@
Discussion. Reformulating various video understanding tasks as visual-language alignment offers several key advantages: (i) it eliminates the need for task-specific architectures by unifying all tasks as interactions between visual and textual features, enabling a streamlined and consistent framework; (ii) it facilitates multitask learning, allowing the model to generalize across task boundaries by leveraging shared spatiotemporal representations. This improves data efficiency and reduces training overhead; (iii) the framework is inherently extensible, supporting the incorporation of new tasks by simply defining additional alignment mechanism, without requiring modifications to the underlying video backbone. By leveraging this unified framework, our method provides a flexible and efficient solution for streaming video understanding across varying complexity and granularity, including video semantics, temporal dependencies, and fine-grained spatial reasoning.
@@ 3.2 Video Encoding
Given a video stream $\mathcal{V}^{T}\in\mathbb{R}^{T\times H\times W\times 3}$, it is first transformed into patch embeddings, with spatial and temporal position embeddings added separately, and the resulting video representation $z\in\mathbb{R}^{T\times hw\times d}$ is the input to the video backbone. We adopt a series of Transformer blocks, that employs divided space-time attention [7], with a key modification to make the temporal attention causal.
@@
Causal Temporal Attention. In contrast to the conventional bi-directional temporal attention used in video understanding backbones [3, 7], causal temporal attention models the temporal correlations between frames in a streaming setting, where frames are processed on-the-fly as they arrive, and each frame is restricted to only access features from preceding frames, i.e., future frames remain inaccessible. This constraint is enforced by applying a temporal causal mask $M\in\mathbb{R}^{T\times T}$ within the attention mechanism. At the $l^{\text{th}}$ layer, the causal temporal attention is formulated as:
@@
Outputs. The entire StreamFormer consists of 12 Transformer blocks, and outputs features of various resolutions: (1) spatiotemporal video features $\mathbf{F}\in\mathbb{R}^{T\times h\times w\times d}$, (2) temporal feature $\mathbf{f}\in\mathbb{R}^{T\times d}$ by applying an attention pooling layer to $\mathbf{F}$ over the spatial dimension following SigLIP [114]. (3) global video feature $\mathbf{v}\in\mathbb{R}^{d}$ by taking the last frame feature of $\mathbf{f}$. All these visual features are projected into a shared visual-text representation space using a visual projector $\phi_{\text{visual}}(\cdot)$. To avoid notation abuse, we reuse $\mathbf{v},\mathbf{f}$ and $\mathbf{F}$ to represent projected features.
@@ 3.3 Language Encoding
As in the former case, we construct prompt templates to convert the category labels into natural language descriptions, e.g., a video clip of kicking a football. In this way, we can reformulate all groundtruth annotations into free-form languages ($y$), and compute the text embeddings ($\textbf{t}\in\mathbb{R}^{d}$) as:
$\textbf{t}=f_{\text{text}}(y),$ (6)
where $f_{\text{text}}(\cdot)$ denotes the pre-trained text encoder from SigLIP. This visual-text alignment facilitates knowledge transfer across tasks with varying granularities, enabling compatibility with diverse label spaces.
@@ 3.4 Spatiotemporal Multitasking
As for training StreamFormer, we unify a wide range of spatiotemporal tasks under a cohesive multitask learning framework. Unlike existing training paradigms that primarily use global contrastive learning on large-scale video-text pairs, we exploit the publicly available datasets from the vision community and focus on learning fine-grained spatiotemporal representations that capture both temporal dynamics and spatial precision. Specifically, we optimize the backbone with three complementary objectives: (i) capturing global semantics through video-level global supervision, (ii) modeling temporal dynamics with frame-level supervision, and (iii) achieving spatial precision via pixel-wise supervision.
@@
$\mathcal{L}_{\text{AR}}=-\frac{1}{|\mathcal{B}|}\sum_{i=1}^{|\mathcal{B}|}\sum_{c=1}^{|\mathcal{C}|}\log\underbrace{\frac{1}{1+e^{y_{ic}(-\tau\mathbf{v}_{i}\cdot\mathbf{t}_{c}+b)}}}_{\mathcal{L}_{ic}},$ (7)
where $y_{ic}$ is the label for video $\mathbf{v}_{i}$ and text $\mathbf{t}_{j}$, which equals 1 if aligned and -1 otherwise. A learnable temperature $\tau$ and bias $b$ are used to alleviate initial imbalance.
For the video-text retrieval loss $\mathcal{L}_{\text{VTR}}$, we simply replace the prompted action label with free-form language description, and calculate the sigmoid loss with negative pairs sampled from the mini-batch [114].
@@
For temporal video grounding, we also formulate it as a contrastive learning paradigm. Specifically, when certain frame in the video falls within the interval corresponding to its paired text query, we consider the pair as a positive sample. Negative samples include: (1) the current frame paired with other language samples in the batch; (2) background frames paired with all language samples in the batch. We calculate $\mathcal{L}_{\text{TVG}}$ following Eq. 8, by replacing the action label with free-form query.
Spatial-level training. Moving to the finest granularity of video understanding, our spatial-level pre-training tackles pixel-wise comprehension through two sub-tasks: video instance segmentation (VIS) for closed-set categories and referring video object segmentation (RVOS) for open-set language queries. VIS aims to predict a class label (including the background) for each pixel in the video, while RVOS needs pixel-level grounding given a natural language query.
To maintain consistency of multitask learning, we directly compute dot product between each patch embeddings with the corresponding textual embeddings to get a coarse-grained segmentation map with logits. After up-sampling the patch-level logits to a higher resolution (e.g. 224$\times$224), a cross-entropy loss is utilized to optimize our predictions to the annotated masks. Formally, the spatial-level loss takes the spatiotemporal feature $\mathbf{F}_{i}=[\mathbf{F}_{i}^{1},\mathbf{F}_{i}^{2},...,\mathbf{F}_{i}^{THW}]\in\mathbb{R}^{(T\times h\times w)\times d}$, where $\mathbf{F}_{i}^{st}\in\mathbb{R}^{d}$ refers to the pixel-level representation at timestep $t$ and spatial location $s$. For closed-set video instance segmentation with classes $\mathcal{C}$, the loss is defined as:
$\mathcal{L}_{\text{VIS}}=-\frac{1}{|\mathcal{B}|THW}\sum_{i=1}^{|\mathcal{B}|}\sum_{st=1}^{THW}\sum_{c=1}^{|\mathcal{C}|}\log\frac{1}{1+e^{y_{ic}^{st}(-\tau\mathbf{F}_{i}^{st}\cdot\mathbf{t}_{c}+b)}},$ (9)
@@
$\mathcal{L}=\underbrace{(\mathcal{L}_{\text{AR}}+\mathcal{L}_{\text{VTR}})}_{\text{global-level}}+\underbrace{(\mathcal{L}_{\text{TAL}}+\mathcal{L}_{\text{TVG}})}_{\text{temporal-level}}+\underbrace{(\mathcal{L}_{\text{VIS}}+\mathcal{L}_{\text{RVOS}})}_{\text{spatial-level}}$
At training time, we alternately sample training data from each task for a forward pass, and use gradient accumulation to perform backpropagation and parameter update collectively after iterating through all tasks, avoiding potential bias towards a specific task. Joint optimization of these objectives allows our model to learn unified video representations capable of capturing global semantics, temporal dynamics and spatial details, all within a single backbone.

### v2

@@ Abstract
Understanding continuous video streams plays a fundamental role in real-time applications, including embodied AI and autonomous driving. Unlike offline video processing, streaming video understanding requires the ability to process video streams frame by frame, preserve historical information, and make low-latency decisions. To address these challenges, our main contributions are three-fold. (i) we develop a novel streaming video backbone, termed as StreamFormer, by incorporating causal temporal attention into a pre-trained vision transformer. This enables efficient streaming video processing while maintaining image representation capability. (ii) to train StreamFormer, we propose to unify diverse spatial-temporal video understanding tasks within a multitask visual-language alignment framework. Hence, StreamFormer learns global semantics, temporal dynamics, and fine-grained spatial relationships simultaneously. (iii) we conduct extensive experiments for online action detection, online video instance segmentation, and video question answering. StreamFormer achieves competitive performance while maintaining efficiency, demonstrating its potential for real-time applications.
@@ 1 Introduction
Comparing to traditional offline approaches, streaming video understanding poses several unique challenges: First, efficient frame-by-frame processing is crucial, as incoming visual signals must be handled promptly and continuously. Unlike existing offline video backbones [35, 3, 102], which typically process clips as their basic unit, streaming models must operate seamlessly at a per-frame basis. Second, long-term context preservation is essential, requiring models to effectively capture temporal dependencies across frames under strict memory and computational constraints. Third, low-latency decision-making is indispensable, ensuring predictions or actions to be promptly delivered for real-time scenarios. Addressing these challenges necessitates specialized backbone architectures specifically designed to balance efficiency, memory management, and accuracy.
@@
Recent works adopt pre-trained image encoders to extract frame features [124, 99], which are subsequently processed auto-regressively by large language models (LLMs). Although this strategy aligns well with streaming video requirements, the reliance on aggregating coarse historical information within LLMs often leads to the loss of important fine-grained spatial and temporal details. To address these limitations, we propose training a streaming video backbone capable of supporting diverse spatiotemporal tasks, ranging from global scene understanding to fine-grained spatiotemporal event recognition. The proposed model, termed as StreamFormer, adapts Vision Transformer architectures pre-trained on images (e.g., SigLIP [119]) by introducing causal attention along the temporal dimension at the spatial-temporal feature extraction stage. This approach allows the backbone to explicitly model temporal dependencies at early vision encoding stages, resulting in richer and more temporally coherent feature representations.
@@
While training the streaming video backbone, we leverage various datasets that are manually annotated by the vision community [12, 38, 107, 11, 2, 37, 113], unifying various spatial-temporal video understanding tasks into a coherent multi-task learning framework. In particular, we combine three complementary learning objectives: (i) video-level classification for global semantics; (ii) per-frame supervision for temporal understanding, and (iii) pixel-wise segmentation for fine-grained spatial comprehension. By leveraging shared textual embeddings from SigLIP, our framework supports both closed-set and open-set tasks. This text-visual alignment mechanism facilitates knowledge transfer across tasks with varying granularities, enabling compatibility with diverse label spaces. In contrast to existing video backbone training paradigms [102, 127, 61] that predominantly rely on global contrastive learning using massive video-text datasets [103, 6, 70], our joint optimization of diverse objectives explicitly ensures capturing detailed temporal motion cues and spatial precision within streaming contexts.
@@
In summary, our contributions are threefold: (i) we propose StreamFormer, which effectively converts pre-trained image backbones into streaming video models via temporal causal attention and spatial low-rank adaptations. Our approach achieves efficient streaming video processing without sacrificing spatial representational quality; (ii) we introduce a hierarchical multitask learning paradigm, systematically unifying global-level, temporal-level, and spatial-level learning objectives. This paradigm enables our model to simultaneously comprehend global semantics, temporal dynamics, and fine-grained spatial relationships within streaming video; (iii) we empirically demonstrate the effectiveness and efficiency of our method on diverse downstream tasks, including online action detection, online video instance segmentation, and video question answering. Our approach achieves promising performance across these tasks, effectively balancing accuracy and computational efficiency in streaming scenarios.
@@ 2 Related Work
Video Backbone. Early video understanding methods employ 3D convolutional neural networks [12, 91, 34] or factorized convolutions [97, 92, 93] to learn spatial-temporal features. Recent works have adopted Vision Transformers [27] to capture long-range dependencies in videos by incorporating global spatiotemporal self-attention [3], local self-attention [68], local-global self-attention [54], divided space-time attention [7], and multi-scale attention mechanisms [33, 57]. Despite the progress in offline video tasks, the use of bi-directional attention in visual backbones makes them less suitable for real-time video tasks. In NLP, LLMs typically adopt a decoder-only architecture [89, 90, 62, 4], where causal attention enables efficient autoregressive generation during inference. Currently, autoregressive modeling in vision primarily focuses on training backbones for image or video generation [94, 75, 5, 85, 87]. For visual understanding tasks, iGPT [18] learns visual representations by predicting next pixels, while AIM [30] explores the scaling laws for autoregressive image pre-training. Toto [81] trains a causal video transformer to predict the next visual patch. Unlike these works, we target designing StreamFormer with causal temporal attention to address various video understanding tasks.
@@
Video Language Pre-training. Building on the success of image-text pretraining [78, 119, 41], recent video-language models [102, 104, 127, 111, 130, 108, 1] learn generalizable representations by training on massive video-text pairs [6, 103, 118]. UMT [53] and InternVideo2 [104] combine feature distilling, masked video modeling and video-text contrastive learning, to learn unimodal and multimodal representations. VideoPrism [127] integrates global-local distillation and token shuffling to enhance video representation learning. Despite strong performance, these methods typically require extensive training on million-scale video-text pairs. In contrast, we explore leveraging the rich, fine-grained human annotations available in existing video datasets for video language training by adopting a multitask learning paradigm, achieving higher data efficiency.
@@
Streaming Video Understanding. Streaming video understanding tasks necessitate the processing of video inputs on a frame-by-frame basis and demand timely reactions [40, 45]. Existing streaming video approaches mainly target at online action detection [109, 110, 16, 96, 128], online video instance segmentation [105, 123, 116] and dense video captioning [133]. Leveraging the autoregressive paradigm in LLMs, recent approaches extract frame-wise features using pre-trained image encoders and pass them to LLMs for online question answering [25, 121, 46]. However, aggregating coarse frame features often results in the loss of fine-grained spatial and temporal details. In contrast, StreamFormer addresses causal temporal interaction in the video backbone to preserve spatial-temporal details.
@@
Multitask Learning. Multitask models provide different outputs for a given input, simultaneously supporting various tasks [13, 83, 117, 29, 71, 49, 15]. Motivated by the success of unifying diverse tasks as sequence-to-sequence modeling in LLMs [80, 9, 23], recent works either train a unified multimodal and multitask transformer model [120, 98, 69, 86, 134, 19], or leverage LLM as a bridge to combine task-specific heads [100, 106, 112]. Inspired by these works, we train our streaming video backbone in a unified multitask visual-language alignment framework.
@@
Figure 2: Overall framework of StreamFormer. Our StreamFormer is trained under a unified visual-language alignment framework, enabling simultaneous understanding of global semantics, temporal dynamics, and fine-grained spatial relationships. Each level utilizes features of different granularities: (i) last frame for the global level, (ii) per frame for the temporal level, (iii) and per frame per patch feature for the spatial level.
@@ 3.1 Problem Formulation
This paper considers the task of streaming video understanding, which aims to continuously process video frames by extracting spatiotemporal features in real-time, while retaining historical context and supporting diverse video understanding tasks. To achieve this, we propose to train a unified streaming video backbone capable of capturing global semantics, temporal dependencies, and fine-grained spatial details, all within a single framework.
Specifically, given a streaming video, represented as $\mathcal{V}^{T}=[I_{1},I_{2},\dots,I_{T}]$, where $T$ denotes the total number of frames in the video. At any time step $t\ (1\leq t\leq T)$, the visual features of different granularities can be computed as:
@@
Discussion. Reformulating various video understanding tasks as visual-language alignment offers several key advantages: (i) it eliminates the need for task-specific architectures by unifying all tasks as interactions between visual and textual features, enabling a streamlined and consistent framework; (ii) it facilitates multitask learning, allowing the model to generalize across task boundaries by leveraging shared spatiotemporal representations. This improves data efficiency and reduces training overhead; (iii) the framework is inherently extensible, supporting the incorporation of new tasks by simply defining additional alignment mechanisms, without requiring modifications to the underlying video backbone. By leveraging this unified framework, our method provides a flexible and efficient solution for streaming video understanding across varying complexity and granularity, including video semantics, temporal dependencies, and fine-grained spatial reasoning.
@@ 3.2 Video Encoding
For the input video $\mathcal{V}^{T}\in\mathbb{R}^{T\times H\times W\times 3}$, it is first transformed into patch embeddings, with spatial and temporal position embeddings added separately, and the resulting video representation $z\in\mathbb{R}^{T\times hw\times d}$ is the input to the video backbone. We adopt a series of Transformer blocks that employ divided space-time attention [7], with a key modification to make the temporal attention causal.
@@
Temporal Causal Attention. In contrast to the conventional bi-directional temporal attention used in video understanding backbones [3, 7], causal temporal attention models the temporal correlations between frames in a streaming setting, where frames are processed on-the-fly as they arrive, and each frame is restricted to only access features from preceding frames, i.e., future frames remain inaccessible. This constraint is enforced by applying a temporal causal mask $M\in\mathbb{R}^{T\times T}$ within the attention mechanism. At the $l^{\text{th}}$ layer, the causal temporal attention is formulated as:
@@
Outputs. StreamFormer consists of 12 Transformer blocks, and outputs features of various resolutions: (1) spatiotemporal video features $\mathbf{F}\in\mathbb{R}^{T\times h\times w\times d}$ by taking the last layer’s hidden states, (2) temporal feature $\mathbf{f}\in\mathbb{R}^{T\times d}$ by applying an attention pooling layer to $\mathbf{F}$ over the spatial dimension following SigLIP [119], (3) global video feature $\mathbf{v}\in\mathbb{R}^{d}$ by taking the last frame feature. All visual features are projected into a shared visual-text representation space using a visual projector $\phi_{\text{visual}}(\cdot)$. To avoid notation abuse, we reuse $\mathbf{v},\mathbf{f}$ and $\mathbf{F}$ to represent projected features.
@@ 3.3 Language Encoding
As in the former case, we construct prompt templates to convert the category labels into natural language descriptions, e.g., a video clip of kicking a football. Details on the prompt templates can be found in the supplementary material. In this way, we can reformulate all groundtruth annotations into free-form languages ($y$), and compute the text embeddings ($\textbf{t}\in\mathbb{R}^{d}$), i.e., $\textbf{t}=f_{\text{text}}(y)$, where $f_{\text{text}}(\cdot)$ denotes the pre-trained text encoder from SigLIP. This visual-text alignment facilitates knowledge transfer across tasks with varying granularities, enabling compatibility with diverse label spaces.
@@ 3.4 Spatiotemporal Multitasking
As for training StreamFormer, we unify a wide range of spatiotemporal tasks under a cohesive multitask learning framework. Unlike existing training paradigms that primarily use global contrastive learning on large-scale video-text pairs, we exploit the publicly available datasets from the vision community and focus on learning fine-grained spatiotemporal representations that capture both temporal dynamics and spatial precision. Specifically, we optimize the backbone with three complementary objectives: (i) capturing global semantics through video-level global supervision, (ii) modeling temporal dynamics with frame-level supervision, and (iii) enabling spatial precision via pixel-wise supervision.
@@
$\mathcal{L}_{\text{AR}}=-\frac{1}{|\mathcal{B}|}\sum_{i=1}^{|\mathcal{B}|}\sum_{c=1}^{|\mathcal{C}|}\log\frac{1}{1+e^{y_{ic}(-\tau\mathbf{v}_{i}\cdot\mathbf{t}_{c}+b)}}$ (6)
where $y_{ic}$ is the label for video $\mathbf{v}_{i}$ and text $\mathbf{t}_{c}$, which equals 1 if aligned and -1 otherwise. A learnable temperature $\tau$ and bias $b$ are used to alleviate initial imbalance.
For the video-text retrieval loss $\mathcal{L}_{\text{VTR}}$, we simply replace the prompted action label with a free-form language description, and calculate the sigmoid loss with negative pairs sampled from the mini-batch [119].
@@
For temporal video grounding, we also formulate it as a contrastive learning paradigm. Specifically, when a certain frame in the video falls within the interval corresponding to its paired text query, we consider the pair as a positive sample. Negative samples include: (1) the current frame paired with other language samples in the batch; (2) background frames paired with all language samples in the batch. We calculate $\mathcal{L}_{\text{TVG}}$ following Eq. 7, by replacing the action label with a free-form query.
Spatial-level training. Moving to the finest granularity, our spatial-level pre-training tackles pixel-wise comprehension through two sub-tasks: video object segmentation (VOS) for closed-set categories and referring video object segmentation (RVOS) for open-set language queries. VOS aims to predict a class label (including the background) for each pixel in the video, while RVOS needs pixel-level grounding given a natural language query.
To maintain the consistency of multitask learning, we directly compute dot product between each patch embedding with the corresponding textual embedding to get a coarse-grained segmentation map with logits. After up-sampling the patch-level logits to a higher resolution (e.g. 224$\times$224), a cross-entropy loss is utilized to optimize our predictions to the annotated masks. Formally, the spatial-level loss takes the spatiotemporal feature $\mathbf{F}_{i}=[\mathbf{F}_{i}^{1},\mathbf{F}_{i}^{2},...,\mathbf{F}_{i}^{THW}]\in\mathbb{R}^{(T\times H\times W)\times d}$, where $\mathbf{F}_{i}^{st}\in\mathbb{R}^{d}$ refers to the pixel-level representation at time $t$ and spatial location $s$. For closed-set video object segmentation with classes $\mathcal{C}$, the loss is defined as:
$\mathcal{L}_{\text{VOS}}=-\frac{1}{|\mathcal{B}|THW}\sum_{i=1}^{|\mathcal{B}|}\sum_{st=1}^{THW}\sum_{c=1}^{|\mathcal{C}|}\log\frac{1}{1+e^{y_{ic}^{st}(-\tau\mathbf{F}_{i}^{st}\cdot\mathbf{t}_{c}+b)}},$ (8)
@@
$\mathcal{L}=\underbrace{(\mathcal{L}_{\text{AR}}+\mathcal{L}_{\text{VTR}})}_{\text{global-level}}+\underbrace{(\mathcal{L}_{\text{TAL}}+\mathcal{L}_{\text{TVG}})}_{\text{temporal-level}}+\underbrace{(\mathcal{L}_{\text{VOS}}+\mathcal{L}_{\text{RVOS}})}_{\text{spatial-level}}$
At training time, we alternately sample data from each task for a forward pass, and use gradient accumulation to perform backpropagation and parameter update collectively after iterating through all tasks, avoiding potential bias towards a specific task. Joint optimization of these objectives allows our model to learn unified video representations capable of capturing global semantics, temporal dynamics and spatial details, all within a single backbone.